OpenStudy (anonymous):

What is the value of 3x2 - 5xy + 5y2 for x = -6 and y = 2? -68 -4 188 208

OpenStudy (anonymous):

=−5xy+5y2+6

OpenStudy (anonymous):

you just have to subtitude the value of x which is -6 in x and 2 in y

OpenStudy (anonymous):

plug in -6 for x and 2 in y

OpenStudy (anonymous):

3(-6)^2 - 5(-6)(2) + 5(2)^2=

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

now just multiply and add and all that fun stuff :)

OpenStudy (anonymous):

3(36) - 5(-12) + 5(4)=

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

108--60 + 20 don't forget that -*-=+ or -- equals +

OpenStudy (anonymous):

108 + 60 +20

OpenStudy (anonymous):

188

OpenStudy (anonymous):

hope it helps

OpenStudy (anonymous):

yes. it did help alot :)